OpenStudy (anonymous):

Find the slope of the line described by 2x + 5y = 10.

OpenStudy (karatechopper):

put it in the form of y=mx+b first..

OpenStudy (karatechopper):

kat i got this lemme do it plz

OpenStudy (anonymous):

-2/5

OpenStudy (karatechopper):

:/

OpenStudy (anonymous):

Thank you :)

OpenStudy (precal):

@karatechopper 2x+5y=10 5y=10-2x y=2-(2/5)x

OpenStudy (karatechopper):

what about em?

OpenStudy (karatechopper):

slope?

OpenStudy (precal):

m=-2/5

OpenStudy (karatechopper):

ok

OpenStudy (karatechopper):

second one u will also put in y=mx+b form